OpenStudy (anonymous):

an automobile with an initial speed of 4.3 m/s accelerates uniformly at the rate of 3.0 m/s^2. A.) What is the final speed of the automobile after 5 s?

OpenStudy (amistre64):

well, considering: s(t) = 1/2 a t^2 + vo sin(alpha) t + so gives us position the rate of change in position defines velocity s'(t) = v(t) = at + vo sin(alpha) for simplicy sin(alpha) = 1 giving us: final velocity = at + initial velocity for a constant acceleration, a